find the smallest square number divisible by each one of the numbers 8,9,10
Answer by amarjeeth123(570)

L.C.M. of 8,9,10 is 2*2*2*3*3*5=360
Since 2 and 5 are not in pairs multiply by 2 and 5 to make it a perfect square.
The required number=360*2*5=3600
The smallest square number divisible by 8,9,10=3600.